# The Nonlinear Filter Model of Stream Cipher Redivivus

### v6

@@ Abstract
The nonlinear filter model is an old and well understood approach to the design of stream ciphers. Extensive research over several decades has shown how to attack stream ciphers based on this model and has identified the security properties required of the Boolean function used as the filtering function to resist such attacks. This led to the problem of constructing Boolean functions which provide adequate security and at the same time are efficient to implement. Unfortunately, over the last two decades no good solutions to this problem appeared in the literature. The lack of good solutions has effectively led to nonlinear filter model becoming more or less obsolete. This is a big loss to the cryptographic design toolkit, since the great advantage of the nonlinear filter model lies, beyond its simplicity and its ability to provide low-cost solutions for hardware-oriented stream ciphers, in the accumulated knowledge about the security requirements for the filtering function, which gives confidence in its security when all criteria are met. In this paper we construct balanced functions on an odd number $n\geq 5$ of variables with the following provable properties: linear bias equal to $2^{-\lfloor n/2\rfloor-1}$, algebraic degree equal to $2^{\lfloor\log_{2}\lfloor n/2\rfloor\rfloor}$, algebraic immunity at least $\lceil(n-1)/4\rceil$, fast algebraic immunity at least $1+\lceil(n-1)/4\rceil$, and the functions can be implemented using $O(n)$ NAND gates. The functions are obtained from a simple modification of the well known class of Maiorana-McFarland bent functions. Due to the efficiency of implementation, for any target security level, we can construct functions which provide the required level of resistance to fast algebraic and fast correlation attacks. By appropriately choosing $n$ and the length $L$ of the linear feedback shift register, we show that it is possible to obtain examples of stream ciphers which are provably $\kappa$-bit secure against well known classes of attacks for various values of $\kappa$. We provide concrete proposals for $\kappa=80,128,160,192,224$ and $256$ using LFSRs of lengths 163, 257, 331, 389, 449, 521 and filtering functions on 75, 119, 143, 175, 203 and 231 variables. For the $80$-bit, $128$-bit, and the $256$-bit security levels, the circuits for the corresponding stream ciphers require about 1743.5, 2771.5, and 5607.5 NAND gates respectively. For the $80$-bit and the $128$-bit security levels, the gate count estimates compare quite well to the famous ciphers Trivium and Grain-128a respectively, while for the $256$-bit security level, we do not know of any other stream cipher design which has such a low gate count.
@@ 1 Introduction
The nonlinear filter model of stream ciphers is several decades old; one may note that the model was extensively discussed in the book by Rueppel [49] which was published in the mid-1980s. The nonlinear filter model consists of two components, namely a linear feedback shift register (LFSR) and a Boolean function $f$ which is applied to a subset of the bits of the LFSR at fixed positions (called tap positions). At each clock, $f$ is applied to the present state of the LFSR to produce a single keystream bit and simultaneously the LFSR also moves to the next state. LFSRs are efficient to implement in hardware. So the implementation efficiency of the nonlinear filter model is essentially determined by the efficiency of implementation of $f$.
@@
Extensive research carried out over the last few decades has shown several approaches to cryptanalysing the nonlinear filter model of stream ciphers. The initial line of attack was based upon determining the linear complexity of the produced keystream. For an LFSR of length $L$ and a Boolean function of algebraic degree $d$, under certain reasonable and easy-to-ensure conditions, the linear complexity of the keystream is known to be at least (${L\choose d}$ (see [49]). Using large enough values of $L$ and $d$, linear complexity based attacks can be made infeasible. The second phase of attacks consisted of various kinds of (fast) correlation attacks. Starting with the first such attack in 1985 [53] (which was efficient on another model, the nonlinear combiner model), a long line of papers [52, 45, 29, 3, 21, 35, 36, 19, 13, 37, 12, 38, 20, 54, 60, 59, 39, 40] explored various avenues for mounting correlation attacks. Surveys of some of the older attacks appear in [10, 11, 43, 2]. Fast correlation attacks are based on affine approximation and apply to the nonlinear filter (as well as the nonlinear combiner) model. The resistance to fast correlation attacks is mainly determined by the linear bias $\varepsilon$ of the Boolean function $f$. The linear bias is determined by the nonlinearity of the function $f$; the higher the nonlinearity, the lower the linear bias. The third phase of attacks started in 2003 with the publication of the algebraic attack [23] and was soon followed by the publication of the fast algebraic attack [22]. Resistance to these attacks requires the function $f$ to possess high (fast) algebraic immunity.
@@
The various attacks mentioned above have posed the following design challenge for a Boolean function to be used in the nonlinear filter model of stream ciphers. Construct balanced Boolean functions which achieve a good combination of nonlinearity and algebraic resistance and are also very efficient to implement. Unfortunately, since the time algebraic attacks were proposed in the early-2000s, no good solutions to the design challenge for Boolean functions have appeared in the literature (some Boolean functions satisfy all the necessary cryptographic criteria, but are too heavy and slow to compute [17], and some others are fast to compute but possess insufficient nonlinearity [55, 57, 15, 42]). A consequence of not being able to find good solutions to the design challenge is that the nonlinear filter model of stream ciphers became obsolete. This is somewhat unfortunate since the model is very old, well studied with well understood security, and the potential to provide low gate count solutions in hardware.
@@
We note that even though the nonlinear filter model became obsolete, use of LFSRs in the design of stream cipher has continued for both hardware and software oriented proposals (see for example [32, 5, 26]). Instead of using a Boolean function, such designs typically use a nonlinear finite state machine to filter the output of the LFSR. Some ciphers such as [9, 4] have gone further and replaced the LFSR with one or more nonlinear feedback shift registers (NFSRs).
@@
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to map $(\mathbf{X},\mathbf{Y})$ to $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of both $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $h$. Assume that $\pi$ is chosen to be an affine map, i.e. the coordinate functions of $\pi$ are affine functions of its input variables. It is known [24] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates. So the obtained bent function has maximum nonlinearity, sufficient (fast) algebraic immunity when the number of variables is large enough (to protect against fast algebraic attacks at a specified security level), and at the same time is quite efficient to implement. Indeed, the (fast) algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation has escaped the notice of earlier researchers. To the best of our knowledge, our result is the first construction of bent functions with a provable lower bound on algebraic immunity.
@@
One problem is that a bent function is not balanced. This problem is easily rectified by XORing a new variable to the bent function, obtaining a function on an odd number of variables. This modification requires only one extra XOR gate for implementation. In terms of security, the modification does not change the linear bias. We prove that the algebraic immunity of the modified function is at least the algebraic immunity of the bent function, and hence the fast algebraic immunity is at least one plus the algebraic immunity of the bent function. So the algebraic resistance of the modified function is essentially the same as that of the bent function. A consequence of unbalancedness of the filtering function is that the keystream sequence is unbalanced, which leads to a distinguishing attack. The simple modification of XORing an extra variable prevents such a distinguishing attack. The resistance to fast correlation attacks and fast algebraic attacks are ensured by the bent function (on an appropriate number of variables). A positive aspect of XORing a new variable is that it prevents (see Theorem 2 of [27]) certain kinds of information leakage which was identified in [3]. On the negative side, filtering functions of the form $W+g(\mathbf{Z})$ were shown to be susceptible to the inversion attack [27]. We prevent this attack by choosing the gap between the first and the last tap positions to be sufficiently large; this countermeasure was already proposed in [27]. More generally, we show that due to the fact that we use filtering functions on a large number of variables, the various kinds of state guessing attacks [28, 33, 47, 58] do not apply to our proposals.
@@
We perform a detailed concrete security analysis of some of the well known attacks on the nonlinear filter model. As the outcome of this analysis, for various security levels, we provide concrete proposals for stream ciphers based on the nonlinear filter model using the Boolean functions described above as the filtering functions. A strong point in favour of these proposals is that at the appropriate security levels they provide provable assurance against well known classes of attacks. Further, we provide concrete gate count estimates for the entire circuit to implement the stream ciphers. For the 80-bit, 128-bit, 160-bit, 192-bit, 224-bit and the $256$-bit security levels, we propose using LFSRs of lengths 163, 257, 331, 389, 449, 521 and filtering functions on 75, 119, 143, 175, 203 and 231 variables respectively. The gate count estimates for the 80-bit, 128-bit, 160-bit, 192-bit, 224-bit and the $256$-bit security levels are 1743.5, 2771.5, 3520.5, 4188.5, 4854.5, and 5607.5 NAND gates respectively. The gate count estimates for the 80-bit and the 128-bit security levels compare quite well [^1] Following [1] we estimated 8 NAND gates for a flip-flop, whereas Trivium estimated 12 NAND gates for a flip-flop. Using 12 NAND gates for a flip-flop, our proposal at the 80-bit security level requires about 2395.5 NAND gates, whereas Trivium requires about 3488 NAND gates. with famous ciphers such as Trivium [9] and Grain-128a [1] which offer 80-bit and 128-bit security respectively. For the other security levels, we are not aware of other stream ciphers which have such low gate counts. So our revival of the nonlinear filter model of stream ciphers leads to concrete proposals which offer a combination of both provable security against well known classes of attacks at a desired level of security and also low gate count.
@@
Well known stream ciphers such as Trivium [9], Mickey [4], Grain-128a [1], SNOW [26] and Sosemanuk [5], use novel and ingenious ideas. Nonetheless, these are standalone designs. The nonlinear filter model, on the other hand, is a model for stream cipher design. Due to the simplicity of the nonlinear filter model, provable properties of the filtering function can be translated into provable resistance of the stream cipher against well known classes of attacks. In particular, for the proposals that we put forward, the provable linear bias of the filtering function translates to provable protection against a large class of fast correlation attacks, and the provable lower bound on the (fast) algebraic immunity translates to provable resistance against algebraic attacks. Stream ciphers based on either nonlinear finite state machines, or nonlinear feedback shift registers do not enjoy this advantage, i.e. for such stream ciphers it is very hard to obtain provable guarantees against various well known types of attacks. As an example, our proposal at the 128-bit security level generates keystream for which the best linear approximation has provable linear bias of $2^{-60}$, while at the same security level, for Sosemanuk [5] the best known [39] linear approximation has correlation $2^{-20.84}$, and it is not known whether there are approximations with higher correlations. Similarly, the time complexities of fast algebraic attacks against Trivium, Mickey, Grain-128a, SNOW and Sosemanuk are not known. It is believed that these stream ciphers can withstand fast algebraic attacks. In contrast, we are able to prove that at the appropriate security levels, the fast algebraic attack is ineffective against the new nonlinear filter model based stream cipher proposals that we put forward.
One advantage of using LFSRs is that it is possible to provably ensure that the LFSR has a maximum period. For stream ciphers based on NFSRs, such as Trivium and Mickey, such provable assurance is not available. For stream ciphers which use a combination of LFSR and NFSR such as Grain-128a, it is possible to mount a divide-and-conquer attack. For example the attack in [54] on Grain-128a finds the state of the LFSR independently of the state of the NFSR. (In some ways this is reminiscent of the attack by Siegenthaler [53] on the nonlinear combiner model.) So even though Grain-128a uses a 256-bit state, due to the divide-and-conquer strategy the full protection of the large state is not achieved. On the other hand, for the nonlinear filter model, there is no known way to mount a divide-and-conquer attack. Our proposal at the 128-bit security level uses a 257-bit LFSR, and there is knwon no way to estimate half the state of the LFSR without involving the other half.
@@
An earlier version of the proposal was attacked [6] using a differential attack. In response, we have modified the proposal to resist the attack in [6]. Though the modified proposal described in this version of the paper resists the attack in [6], we do not have any proof that the present proposal resists all kinds of differential attacks. We note that while provable properties of the filtering function can be translated to provable resistance against certain classes of attacks, it is by no means true that these properties provide resistance against all classes of attacks. We welcome further analysis and investigation of other avenues of attacks on the concrete stream cipher proposals that we put forward in this paper.

### v7

@@ Abstract
The nonlinear filter model is an old and well understood approach to the design of secure stream ciphers. Extensive research over several decades has shown how to attack stream ciphers based on this model and has identified the security properties required of the Boolean function used as the filtering function to resist such attacks. This led to the problem of constructing Boolean functions which provide adequate security and at the same time are efficient to implement. Unfortunately, over the last two decades no fully satisfactory solutions to this problem appeared in the literature. The lack of good solutions has effectively led to the nonlinear filter model becoming more or less obsolete. This is a big loss to the cryptographic design toolkit, since the great advantages of the nonlinear filter model are its simplicity, well understood security and the potential to provide low cost solutions for hardware oriented stream ciphers. In this paper, we revive the nonlinear filter model by constructing appropriate Boolean functions which provide required security and are also efficient to implement. We put forward concrete suggestions of stream ciphers which are $\kappa$-bit secure against known types of attacks for $\kappa=80$, 128, 160, 192, 224 and 256. For the 80-bit and the 128-bit security levels, the gate count estimates of our proposals compare quite well to the famous ciphers Trivium and Grain-128a respectively, while for the 256-bit security level, we do not know of any other stream cipher design which has such a low gate count.
@@ 1 Introduction
The nonlinear filter model of stream ciphers is several decades old; one may note that the model was extensively discussed in the book by Rueppel [49] which was published in the mid-1980s. The nonlinear filter model consists of two components, namely a linear feedback shift register (LFSR) and a Boolean function $f$ which is applied to a subset of the bits of the LFSR at fixed positions (called tap positions). At each clock, $f$ is applied to the present state of the LFSR to produce a single keystream bit and simultaneously the LFSR also moves to the next state. LFSRs are efficient to implement in hardware. So the implementation efficiency of the nonlinear filter model is essentially determined by the efficiency of implementing $f$.
@@
A basic requirement on the filtering function $f$ is that it is balanced, as otherwise the keystream sequence produced by the stream cipher is unbalanced, which leads to a distinguishing attack. Extensive research carried out over the last few decades has shown several approaches to cryptanalysing the nonlinear filter model of stream ciphers. The initial line of attack was based upon determining the linear complexity of the produced keystream. For an LFSR of length $L$ and a Boolean function of algebraic degree $d$, under certain reasonable and easy-to-ensure conditions, the linear complexity of the keystream is known to be at least $L\choose d$} (see [49]). Using large enough values of $L$ and $d$, linear complexity based attacks can be made infeasible. The second phase of attacks consisted of various kinds of (fast) correlation attacks. Starting with the first such attack in 1985 [53] (which was efficient on another model, the nonlinear combiner model), a long line of papers [52, 45, 29, 3, 21, 34, 35, 19, 13, 36, 12, 37, 20, 54, 60, 59, 39, 40] explored various avenues for mounting correlation attacks. Surveys of some of the older attacks appear in [10, 11, 43, 2]. Fast correlation attacks are based on affine approximation and apply to the nonlinear filter (as well as the nonlinear combiner) model. The resistance to fast correlation attacks is mainly determined by the linear bias of the Boolean function $f$. The linear bias is determined by the nonlinearity of the function $f$; the higher the nonlinearity, the lower the linear bias. The third phase of attacks started in 2003 with the publication of the algebraic attack [23] and was soon followed by the publication of the fast algebraic attack [22]. Resistance to these attacks requires the function $f$ to possess high (fast) algebraic immunity.
@@
The various attacks mentioned above have posed the following design challenge for a Boolean function to be used in the nonlinear filter model of stream ciphers. Construct balanced Boolean functions which achieve a good combination of nonlinearity and algebraic resistance and are also very efficient to implement. Unfortunately, since the time algebraic attacks were proposed in the early-2000s, no good solutions to the design challenge for Boolean functions have appeared in the literature (some Boolean functions satisfy all the necessary cryptographic criteria, but are too heavy and slow to compute [17], and some others are fast to compute but possess insufficient nonlinearity [55, 57, 15, 42]). A consequence of not being able to find good solutions to the design challenge is that the nonlinear filter model of stream ciphers became obsolete. This is somewhat unfortunate since the model is very old, well studied with well understood security, and has the potential to provide low gate count solutions in hardware.
@@
We note that even though the nonlinear filter model became obsolete, use of LFSRs in the design of stream ciphers has continued for both hardware and software oriented proposals (see for example [32, 5, 26]). Instead of using a Boolean function, such designs typically use a nonlinear finite state machine to filter the output of the LFSR. Some ciphers such as [9, 4] have gone further and replaced the LFSR with one or more nonlinear feedback shift registers (NFSRs).
@@
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to map $(\mathbf{X},\mathbf{Y})$ to $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of both $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $h$. Assume that $\pi$ is chosen to be an affine map, i.e. the coordinate functions of $\pi$ are affine functions of its input variables. It is known [24] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates (this may seem a folklore result, but as far as we are aware, this result never appeared earlier in the literature.) So the obtained bent function has maximum nonlinearity, sufficient (fast) algebraic immunity when the number of variables is large enough (to protect against fast algebraic attacks at a specified security level), and at the same time is quite efficient to implement. Indeed, the (fast) algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation has escaped the notice of earlier researchers. To the best of our knowledge, our result is the first construction of bent functions with a provable non-trivial lower bound on algebraic immunity.
@@
One problem is that a bent function is not balanced. This problem is easily rectified by XORing a new variable to the bent function, obtaining a function on an odd number of variables. This modification requires only one extra XOR gate for implementation. In terms of security, the modification does not change the linear bias. We prove that the algebraic immunity of the modified function is at least the algebraic immunity of the bent function, and hence the fast algebraic immunity is at least one plus the algebraic immunity of the bent function. So the algebraic resistance of the modified function is essentially the same as that of the bent function. The resistance to fast correlation attacks and fast algebraic attacks are ensured by the bent function (on an appropriate number of variables). A positive aspect of XORing a new variable is that it prevents (see Theorem 2 of [27]) certain kinds of information leakage which was identified in [3]. On the negative side, filtering functions of the form $W+g(\mathbf{Z})$ were shown to be susceptible to the inversion attack [27]. We prevent this attack by choosing the gap between the first and the last tap positions to be sufficiently large; this countermeasure was already proposed in [27]. More generally, we show that due to the fact that we use filtering functions on a large number of variables, the various kinds of state guessing attacks [28, 33, 47, 58] do not apply to our proposals.
@@
We perform a detailed concrete security analysis of some of the well known attacks on the nonlinear filter model. As the outcome of this analysis, for various security levels, we provide concrete proposals for stream ciphers based on the nonlinear filter model using the Boolean functions described above as the filtering functions. A strong point in favour of these proposals is that at the appropriate security levels they provide provable assurance against well known classes of attacks. Further, we provide concrete gate count estimates for the entire circuit to implement the stream ciphers. For the 80-bit, 128-bit, 160-bit, 192-bit, 224-bit and the $256$-bit security levels, we propose using LFSRs of lengths 163, 257, 331, 389, 449, 521 and filtering functions on 75, 119, 143, 175, 203 and 231 variables respectively. The gate count estimates for the 80-bit, 128-bit, 160-bit, 192-bit, 224-bit and the $256$-bit security levels are 1797.5, 2840.5, 3596.0, 4265.5, 4942.5, and 5693.0 NAND gates respectively. The gate count estimates for the 80-bit and the 128-bit security levels compare quite well [^1] Following [1] we estimated 8 NAND gates for a flip-flop, whereas Trivium estimated 12 NAND gates for a flip-flop. Using 12 NAND gates for a flip-flop, our proposal at the 80-bit security level requires about 2464.5 NAND gates, whereas Trivium requires about 3488 NAND gates. with famous ciphers such as Trivium [9] and Grain-128a [1] which offer 80-bit and 128-bit security respectively. For the other security levels, we are not aware of other stream ciphers which have such low gate counts. So our revival of the nonlinear filter model of stream ciphers leads to concrete proposals which offer a combination of both provable security against well known classes of attacks at a desired level of security and also low gate count.
@@
Well known stream ciphers such as Trivium [9], Mickey [4], Grain-128a [1], SNOW [26] and Sosemanuk [5], use novel and ingenious ideas. Nonetheless, these are standalone designs. The nonlinear filter model, on the other hand, is a model for stream cipher design. Due to the simplicity of the nonlinear filter model, provable properties of the filtering function can be translated into provable resistance of the stream cipher against well known classes of attacks. In particular, for the proposals that we put forward, the provable linear bias of the filtering function translates to provable protection against a large class of fast correlation attacks, and the provable lower bound on the (fast) algebraic immunity translates to provable resistance against algebraic attacks. Stream ciphers based on either nonlinear finite state machines, or nonlinear feedback shift registers do not enjoy this advantage, i.e. for such stream ciphers it is very hard to obtain provable guarantees against various well known types of attacks. As an example, our proposal at the 128-bit security level generates keystream for which the best linear approximation has provable linear bias of $2^{-60}$, while at the same security level, for Sosemanuk [5] the best known [39] linear approximation has correlation $2^{-20.84}$, and it is not known whether there are approximations with higher correlations. Similarly, the time complexities of fast algebraic attacks against Trivium, Mickey, Grain-128a, SNOW and Sosemanuk are not known. It is believed that these stream ciphers can withstand fast algebraic attacks. In contrast, we provide strong security guarantee that at the appropriate security levels, the fast algebraic attack is ineffective against the new nonlinear filter model based stream cipher proposals that we put forward.
One advantage of using LFSRs is that it is possible to provably ensure that the LFSR has a maximum period. For stream ciphers based on NFSRs, such as Trivium and Mickey, such provable assurance is not available. For stream ciphers which use a combination of LFSR and NFSR such as Grain-128a, it is possible to mount a divide-and-conquer attack. For example the attack in [54] on Grain-128a finds the state of the LFSR independently of the state of the NFSR. (In some ways this is reminiscent of the attack by Siegenthaler [53] on the nonlinear combiner model.) So even though Grain-128a uses a 256-bit state, due to the divide-and-conquer strategy the full protection of the large state is not achieved. On the other hand, for the nonlinear filter model, there is no known way to mount a divide-and-conquer attack. Our proposal at the 128-bit security level uses a 257-bit LFSR, and there is no known way to estimate half the state of the LFSR without involving the other half.
@@
An earlier version of the proposal was attacked [6] using a differential attack. In response, we have modified the proposal to resist the attack in [6]. Though the modified proposal described in this version of the paper resists the attack in [6], we do not have any proof that the present proposal resists all kinds of differential attacks. We note that while provable properties of the filtering function can be translated to provable resistance against certain classes of attacks, it is by no means true that these properties provide resistance against all classes of attacks.
Reference implementation.
We welcome further study of the concrete proposals that we have put forward. To help in such study, we have made a reference implementation of the proposals. The simple C code for the reference implementation can be obtained from the following link.
https://github.com/palsarkar/nonlinear-filter-stream-cipher